Hello i ned help solving this equation:
(^2-3)(^2+3)
THANK YOU SOOO MUCH!!

Answer by mmm4444bot(95)

Hello There:
(^2-3)(^2+3)
This is not an equation. An equation always contains an equals sign.
Additionally, these expressions do not make sense because you cannot begin with the ^symbol.
If you mean:
(x^2 - 3)*(x^2 + 3)
and the instructions are to simplify, then this is the factorization of a "difference of squares" -- which should be discussed in your text.
The pattern is:
(a^2 - b^2) = (a + b)*(a - b)
In your problem, a = x^2 and b = 3.
Squaring both and taking their difference gives us:
x^4 - 9
~ Mark